# DAISI: Data Assimilation with Inverse Sampling using Stochastic Interpolants

### v1

@@ A.5.1 Diffusion Posterior Sampling (DPS)
DPS (chung2023diffusion) proceeds by approximating $p({\bm{y}}\mid{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}\mid{\bm{z}}_{1})]$ by simply taking the expectation inside the likelihood:
@@
When $p({\bm{y}}\mid{{\bm{z}}}_{1})$ is a Gaussian, we use the additional scaling factor proposed by (chung2023diffusion) which gives the expression
@@ A.5.2 Moment-Matching Posterior Sampling (MMPS)
MMPS extends DPS by considering the approximation $p({\bm{z}}_{1}|{\bm{z}}_{t})\approx\mathcal{N}(\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}],\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}])$, where $\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ is obtained as before, and $\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}]:=\mathbb{E}[{\bm{z}}_{1}{\bm{z}}_{1}^{\top}|{\bm{z}}_{t}]-\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]^{\top}$ can be computed using the following formula (Rozet_Andry_Lanusse_Louppe_2024):
@@ A.7 Probability flow ODE
Diffusion models work with SDEs instead of ODEs but can still be used within our framework through the so-called Probability Flow ODE (song2020score), which we demonstrate here.
@@ Appendix C Entropy dissipation
In this section, we prove a Bakry-Émery-type entropy dissipation result (Proposition 3.1) (bakry2006diffusions), which describes the influence of the stochastic parameter $\epsilon$ on the information loss of the forecast ensembles when assimilating data with DAISI.
@@
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. That is,
@@
$\mathrm{d}{\bm{z}}_{\tau}=-{\bm{b}}_{1-\tau}({\bm{x}})\mathrm{d}\tau+\epsilon\nabla\log p_{1-\tau}({\bm{z}}_{\tau})+\sqrt{2\epsilon}\,\mathrm{d}W_{\tau},\quad z_{0}\sim\hat{\pi}_{n},$ (96)
@@
$-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}})q_{\tau}({\bm{x}}))\mathrm{d}{\bm{x}}$ $=-\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}}))\mathrm{d}{\bm{x}}-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}}){\bm{b}}_{1-t}({\bm{x}})}{q_{\tau}({\bm{x}})}\nabla q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}$ (106)
@@
$\mathrm{d}{\bm{z}}_{t}={\bm{b}}^{\bm{y}}_{t}({\bm{x}})\mathrm{d}t+\epsilon\nabla\log p^{\bm{y}}_{t}({\bm{z}}_{t})+\sqrt{2\epsilon}\,\mathrm{d}W_{t},\quad z_{0}\sim\hat{\rho}_{0}^{\epsilon},$ (114)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})=-2\lambda\epsilon\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})$ (117)
@@ D.1 Metrics
To measure the calibration of the ensemble forecasts, we measure the Continuous Ranked Probability Score (CRPS) (gneiting2007strictly). We follow and compute the fair unbiased CRPS estimate (fairCRPS; zamo2018estimation) for our ensemble, which reads
@@
Additionally, we measure the Spread Skill Ratio (SSR) to evaluate the ensemble calibration, where a well calibrated ensemble should have $\text{SSR}\approx 1$ (fortin2014should). This is defined as
@@
$\text{Spread}^{t}_{d}=\sqrt{\frac{1}{|G|N_{\text{ens}}}\sum_{g\in G}\sum^{N_{ens}}_{ens=1}(\bar{\hat{X}}^{t,ens}_{g,d}-\hat{X}^{t,ens}_{g,d})^{2}}.$ (127)
@@
When we have access to samples $\{{\bm{x}}_{i}\}_{i=1}^{N}$, $\{{\bm{y}}_{j}\}_{j=1}^{M}$ from two measures $\pi_{1}$ and $\pi_{2}$, respectively, we can compute a particular notion of distance between them, using the Maximum Mean Discrepancy (MMD) (gretton2012kernel), defined as the distance between the kernel embeddings of the two measures in the corresponding reproducing kernel Hilbert space (RKHS) $\mathcal{H}$. In practice, given a kernel $k(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, this is computed as
@@ D.3 Lorenz ’63
The Lorenz ’63 (L63) model (lorenz1963deterministic) is given by the following system of ODEs
@@ D.4 Surface Quasi-Geostrophic (SQG)
For the SQG experiments, we use the two-layer surface quasi-geostrophic model presented in (tulloch2009quasigeostrophic). This formulation evolves potential temperature at both the lower boundary ($z=0$) and an upper lid ($z=H$), capturing the influence of surface and tropopause-level dynamics on the interior flow. The three-dimensional geostrophic velocity field is diagnosed via a nonlocal spectral inversion, which couples the boundaries through stratified Green’s functions. Compared to classical SQG, which assumes decay into the deep interior ($z\to\infty$), the two-layer model confines dynamics to a finite slab and supports richer vertical structure, including baroclinic interactions.
The equations are solved numerically by first applying a fast Fourier transform (FFT) to map model variables to spectral space. They are then integrated forward with a fourth order Runge-Kutta solver that uses a 2/3 dealiasing rule and implicit treatment of hyperdiffusion. For more details, we refer to the GitHub repository of the model (sqgturb).
@@ D.4.1 Model training
Our backbone model for learning the drift $\bm{b}$} is a modified U-Net (UNET_Original) following the design of song2020score and Karras2022edm with 3.5M parameters. Since the SQG data is 2D periodic, we use circular padding to preserve spatial continuity. The same network configuration is used for both the $64\times 64$ and $256\times 256$ experiments, but each model is trained independently. The U-Net employs a hidden dimension of 32 throughout and consists of three hierarchical levels, with attention applied at the second level. Although we did not explore more advanced architectures, our framework is compatible with any model capable of learning a flow-matching objective.
@@
Optimiser AdamW (loshchilov_fixing_2017)
Initialization Xavier Uniform (glorot_understanding_2010)
LR decay schedule Cosine (loshchilov_sgdr_2017)
@@ D.4.2 Data assimilation setup
$y_{n}=\mathcal{H}(x_{n})+\eta_{n},\qquad\mathcal{H}(x_{n})=(A\circ h)(x_{n}),\qquad\eta_{n}\sim\mathcal{N}(0,\sigma_{\mathrm{obs}}^{2}),$ (142)
where $h$ is the operator in Table 2 and $A$ is the sparsity operator. The observation locations in $A$ are chosen uniformly at random before assimilation and remain the same across time and members.
@@
There are several options for initializing DAISI, depending on the available initial conditions. In our numerical experiments, assimilation begins from a noisy version of the true initial state $\tilde{\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0},\sigma_{\text{init}})$. This perturbation is needed for LETKF to work, and is thus also used by DAISI to ensure fairness. In particular, this fairness is important for the squared observations, where the initial conditions affect how multimodal the filtering becomes.
@@
This initialization procedure is reminiscent of SDEdit (meng2022sdedit), in which a sample is partially inverted by the stochastic interpolant ${\bm{z}}_{t}=\alpha_{t}{\bm{z}}_{1}+\beta_{t}{\bm{z}}_{0}$. This has been used for image editing, and in Table 11, we perform an ablation where it replaces the backward SDE in the inversion step.
@@ D.4.3 Baselines
We compare DAISI against both classical and ML-based DA methods. Classical baselines include the Localized Ensemble Transport Filter (LETKF), while ML baselines include FlowDAS and the Ensemble Score Filter (EnSF).
@@
We used the Ensemble Score Filter (EnSF) as described in (liang2025ensemble) with $\epsilon_{\alpha}=0.05$ and 1000 Euler steps. We note that the EnSF implementation includes an inflation step not mentioned in the original article, which we found crucial to prevent mode collapse. After each assimlation cycle, the particles are updated as
@@
We trained a conditional stochastic interpolant model for probabilistic forecasting (chen2024probabilistic), as used in FlowDAS, to emulate the forward dynamics of SQG at $64\times 64$ resolution. Our model conditions on the past six states for roll-out, i.e., generates samples from $p({\bm{x}}_{n}|{\bm{x}}_{n-1},\ldots,{\bm{x}}_{n-6})$. For the drift model $\bm{b}$}, we used a U-Net, similar to the architecture used for the interpolant in DAISI, with $64$ channels, and consisting of three hierarchical levels with channel multipliers $(1,2,2)$. Each level uses group-normalized residual blocks with 8 groups, and time embeddings are provided through learned sinusoidal features of dimension 32. Linear self-attention is applied at every resolution, together with a full multi-head self-attention block applied at the bottleneck between the encoder and decoder. The model is trained using AdamW with learning rate $2\times 10^{-4}$, cosine learning-rate decay, gradient-norm clipping, batch size 32, and dataset normalization with standard deviation $\sigma_{\text{train}}=2660$. Training is performed for 4096 epochs on 3-hourly SQG data.
For sampling, we use the parameters in Table 6, which are based on (chen2025flowdas) and finetuned for the different cases.
@@
LETKF is a state-of-the-art DA method commonly used in the geosciences (wang2021multiscale). The hyperparameters used for LETKF are shown in Table 7. These are roughly based on the ones in (liang2025ensemble) but finetuned for the different cases. For the high-dimensional case, no parameter search is done due to computational cost.

### v2

@@ A.5.1 Diffusion Posterior Sampling (DPS)
DPS (Chung et al., 2023) proceeds by approximating $p({\bm{y}}\mid{\bm{z}}_{t})=\mathbb{E}_{{\bm{z}}_{1}|{\bm{z}}_{t}}[p({\bm{y}}\mid{\bm{z}}_{1})]$ by simply taking the expectation inside the likelihood:
@@
When $p({\bm{y}}\mid{{\bm{z}}}_{1})$ is a Gaussian, we use the additional scaling factor proposed by (Chung et al., 2023) which gives the expression
@@ A.5.2 Moment-Matching Posterior Sampling (MMPS)
MMPS extends DPS by considering the approximation $p({\bm{z}}_{1}|{\bm{z}}_{t})\approx\mathcal{N}(\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}],\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}])$, where $\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]$ is obtained as before, and $\mathbb{V}[{\bm{z}}_{1}|{\bm{z}}_{t}]:=\mathbb{E}[{\bm{z}}_{1}{\bm{z}}_{1}^{\top}|{\bm{z}}_{t}]-\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]\mathbb{E}[{\bm{z}}_{1}|{\bm{z}}_{t}]^{\top}$ can be computed using the following formula (Rozet et al., 2024):
@@ A.7 Probability flow ODE
Diffusion models work with SDEs instead of ODEs but can still be used within our framework through the so-called Probability Flow ODE (Song et al., 2021), which we demonstrate here.
@@ Appendix C Entropy dissipation
In this section, we prove a Bakry-Émery-type entropy dissipation result (Proposition 3.1) (Bakry and Émery, 2006), which describes the influence of the stochastic parameter $\epsilon$ on the information loss of the forecast ensembles when assimilating data with DAISI.
@@
Assume that the marginal laws $\{\rho_{t}\}_{t\in[0,1]}$ and $\{\rho_{t}^{\bm{y}}\}_{t\in[0,1]}$ of the interpolants bridging $\rho_{0}=\mathcal{N}(\bm{0},\mathbf{I})$ with $\rho_{1}=\mathbb{P}_{\infty}$ and $\rho_{1}^{{\bm{y}}}=\mathbb{P}_{\infty}^{\bm{y}}$, respectively, satisfy a log-Sobolev inequality with constant $\lambda>0$, uniform in $t$. That is,
@@
$\mathrm{d}{\bm{z}}_{\tau}=-{\bm{b}}_{1-\tau}({\bm{z}}_{\tau})\mathrm{d}\tau+\epsilon\nabla\log p_{1-\tau}({\bm{z}}_{\tau})+\sqrt{2\epsilon}\,\mathrm{d}W_{\tau},\quad z_{0}\sim\hat{\pi}_{n},$ (97)
@@
$-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}})q_{\tau}({\bm{x}}))\mathrm{d}{\bm{x}}$ $=-\int_{\mathbb{R}^{d_{x}}}r_{\tau}({\bm{x}})\mathrm{div}({\bm{b}}_{1-\tau}({\bm{x}}))\mathrm{d}{\bm{x}}-\int_{\mathbb{R}^{d_{x}}}\frac{r_{\tau}({\bm{x}}){\bm{b}}_{1-\tau}({\bm{x}})}{q_{\tau}({\bm{x}})}\nabla q_{\tau}({\bm{x}})\mathrm{d}{\bm{x}}$ (107)
@@
$\mathrm{d}{\bm{z}}_{t}={\bm{b}}^{\bm{y}}_{t}({\bm{z}}_{t})\mathrm{d}t+\epsilon\nabla\log p^{\bm{y}}_{t}({\bm{z}}_{t})+\sqrt{2\epsilon}\,\mathrm{d}W_{t},\quad z_{0}\sim\hat{\rho}_{0}^{\epsilon},$ (115)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})\leq-2\lambda\epsilon\mathcal{KL}(r_{t}^{\bm{y}}||p_{t}^{\bm{y}})$ (118)
@@ D.1 Metrics
To measure the calibration of the ensemble forecasts, we measure the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007). We follow and compute the fair unbiased CRPS estimate (Ferro, 2014; Zamo and Naveau, 2018) for our ensemble, which reads
@@
Additionally, we measure the Spread Skill Ratio (SSR) to evaluate the ensemble calibration, where a well calibrated ensemble should have $\text{SSR}\approx 1$ (Fortin et al., 2014). This is defined as
@@
$\text{Spread}^{t}_{d}=\sqrt{\frac{1}{|G|N_{\text{ens}}}\sum_{g\in G}\sum^{N_{ens}}_{ens=1}(\bar{\hat{X}}^{t}_{g,d}-\hat{X}^{t,ens}_{g,d})^{2}}.$ (128)
@@
When we have access to samples $\{{\bm{x}}_{i}\}_{i=1}^{N}$, $\{{\bm{y}}_{j}\}_{j=1}^{M}$ from two measures $\pi_{1}$ and $\pi_{2}$, respectively, we can compute a particular notion of distance between them, using the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012), defined as the distance between the kernel embeddings of the two measures in the corresponding reproducing kernel Hilbert space (RKHS) $\mathcal{H}$. In practice, given a kernel $k(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, this is computed as
@@ D.3 Lorenz ’63
The Lorenz ’63 (L63) model (Lorenz, 1963) is given by the following system of ODEs
@@ D.4 Surface Quasi-Geostrophic (SQG)
For the SQG experiments, we use the two-layer surface quasi-geostrophic model presented in (Tulloch and Smith, 2009). This formulation evolves potential temperature at both the lower boundary ($z=0$) and an upper lid ($z=H$), capturing the influence of surface and tropopause-level dynamics on the interior flow. The three-dimensional geostrophic velocity field is diagnosed via a nonlocal spectral inversion, which couples the boundaries through stratified Green’s functions. Compared to classical SQG, which assumes decay into the deep interior ($z\to\infty$), the two-layer model confines dynamics to a finite slab and supports richer vertical structure, including baroclinic interactions.
The equations are solved numerically by first applying a fast Fourier transform (FFT) to map model variables to spectral space. They are then integrated forward with a fourth order Runge-Kutta solver that uses a 2/3 dealiasing rule and implicit treatment of hyperdiffusion. For more details, we refer to the GitHub repository of the model (Whitaker, 2025).
@@ D.4.1 Model training
Our backbone model for learning the drift $\bm{b}$} is a modified U-Net (Ronneberger et al., 2015) following the design of Song et al. (2021) and Karras et al. (2022) with 3.5M parameters. Since the SQG data is 2D periodic, we use circular padding to preserve spatial continuity. The same network configuration is used for both the $64\times 64$ and $256\times 256$ experiments, but each model is trained independently. The U-Net employs a hidden dimension of 32 throughout and consists of three hierarchical levels, with attention applied at the second level. Although we did not explore more advanced architectures, our framework is compatible with any model capable of learning a flow-matching objective.
@@
Optimiser AdamW (Loshchilov and Hutter, 2017a)
Initialization Xavier Uniform (Glorot and Bengio, 2010)
LR decay schedule Cosine (Loshchilov and Hutter, 2017b)
@@ D.4.2 Data assimilation setup
$y_{n}=\mathcal{H}(x_{n})+\eta_{n},\qquad\mathcal{H}(x_{n})=(A\circ h)(x_{n}),\qquad\eta_{n}\sim\mathcal{N}(0,\sigma_{\mathrm{obs}}^{2}I),$ (143)
where $h$ is the operator in Table 2 and $A$ is the sparsity operator. The observation locations in $A$ are chosen uniformly at random before assimilation and remain the same across time and members and channels.
For the High-dim. experiments on $256\times 256$, observations are generated through a three-step process. The state is first subsampled to $64\times 64$, after which each grid point is replaced by the average over a $5\times 5$ neighborhood, mimicking lower-resolution sensing. Finally, only 5% of these pixels are observed at random with added noise, yielding highly sparse observations.
@@
There are several options for initializing DAISI, depending on the available initial conditions. In our numerical experiments, assimilation begins from a noisy version of the true initial state $\tilde{\bm{x}}_{0}\sim\mathcal{N}({\bm{x}}_{0},\sigma_{\text{init}}^{2}\mathbf{I})$. This perturbation is needed for LETKF to work, and is thus also used by DAISI to ensure fairness. In particular, this fairness is important for the squared observations, where the initial conditions affect how multimodal the filtering becomes.
@@
This initialization procedure is reminiscent of SDEdit (Meng et al., 2022), in which a sample is partially inverted by the stochastic interpolant ${\bm{z}}_{t}=\alpha_{t}{\bm{z}}_{1}+\beta_{t}{\bm{z}}_{0}$. This has been used for image editing, and in Table 14, we perform an ablation where it replaces the backward SDE in the inversion step.
@@ D.4.3 Baselines
We compare DAISI against both classical and ML-based DA methods. Classical baselines include the Local Ensemble Transform Kalman Filter (LETKF), while ML baselines include score-based data assimilation (SDA), FlowDAS, and the Ensemble Score Filter (EnSF).
SDA (smoothing):
The score-based data assimilation (SDA) algorithm of (Rozet and Louppe, 2023) uses a diffusion model trained on a short window of a dynamical system’s trajectory, referred to as the Markov blanket, to sample trajectories from the smoothing distribution $p({\bm{x}}_{1:N}|{\bm{y}}_{1:N})$ of an arbitrary length $N$, using a guidance-based method for posterior sampling. To be consistent with the other baselines, we also conditioned on noised initial states $\tilde{{\bm{x}}}_{0}^{(j)}$ as described in Section D.4.2, yielding samples from the distribution $p({\bm{x}}_{0:N}|{\bm{y}}_{1:N},\tilde{{\bm{x}}}_{0}^{(j)})$. For the window length $W$ of the Markov blanket, we set $W=5$ in all of our experiments. For the score network, we used the default U-net architecture found in the original SDA github repository (https://github.com/francois-rozet/sda) and trained the model for 4096 epochs using the AdamW optimizer, with learning rate $10^{-3}$, weight decay $10^{-3}$ and linear training scheduler. We use the optimal hyperparameters in Table 6, obtained by grid search.
We note, however, that SDA is a smoothing algorithm and therefore does not provide a completely fair comparison with DAISI and the other baselines, which are essentially filtering algorithms. For this purpose, we also propose a filtering variant of SDA that we describe in the following.
Table 6: Hyperparameter configurations for SDA smoothing.
Experiment Guidance method Guidance strength $\zeta$ Euler steps Correction steps
Noisy MMPS 1 100 1
Sparse MMPS 1 500 1
Multimodal MMPS 10 250 2
Saturating MMPS 5 100 2
SEVIR MMPS 1 500 2
SDA (filtering):
In order to perform approximate filtering using the spatio-temporal diffusion model trained for SDA with window size $W$, we propose to iteratively sample ${\bm{x}}_{n:n+W-1}^{(j)}\sim p({\bm{x}}_{n:n+W-1}|{\bm{y}}_{n+1:n+W-1},\tilde{{\bm{x}}}_{n}^{(j)})$ for $n=0,\ldots,N-W+1$ and $j=1,\ldots,J$, storing ${\bm{x}}_{n+W-1}^{(j)}$ as the filtered state at time step n+W1n+W-1 and setting $\tilde{{\bm{x}}}_{n+1}^{(j)}\leftarrow{\bm{x}}_{n+1}^{(j)}$ for the “initial condition” in the next iteration. We summarize this in Algorithm 2. We use a window size $W=5$ for all of our experiments and the same the score network as used in the smoothing variant of SDA. The hyperparameters used are displayed in Table 7. We also note that this setup is similar to the joint guided autoregressive model proposed in (Shysheya et al., 2024).
Algorithm 2 SDA filtering
1:Inputs: Initial ensemble $\tilde{x}_{0}^{(j)}\sim\mathcal{N}({\bm{x}}_{0},\sigma_{\text{init}}^{2}\mathbf{I})$ for $j=1,\ldots,J$, observations $\{{\bm{y}}_{n}\}_{n=1}^{N}$, and $A=\varnothing$
2:for $n=0,\dots,N-W+1$ do
3:  Initialize $B_{n}=\varnothing$
4:  for $j=1,\dots,J$ do
5:   Sample ${\bm{x}}_{n:n+W-1}^{(j)}\sim p({\bm{x}}_{n:n+W-1}|{\bm{y}}_{n+1:n+W-1},\tilde{{\bm{x}}}_{n}^{(j)})$ via guidance
6:   Append last state ${\bm{x}}_{n+W-1}^{(j)}$ to $B_{n}$
7:   Set $\tilde{{\bm{x}}}_{n+1}^{(j)}\leftarrow{\bm{x}}_{n+1}^{(j)}$
8:  end for
9:  Append $B_{n}$ to AA
10:end for
11:Output: Sequence of SDA filtered states $A=\{\{{\bm{x}}_{n}^{(j)}\}_{j=1}^{J}\}_{n=W-1}^{N}$
Table 7: Hyperparameter configurations for SDA filtering.
Experiment Guidance method Guidance strength $\zeta$ Euler steps Correction steps
Noisy MMPS 1 100 1
Sparse MMPS 1 100 1
Multimodal MMPS 5 100 1
Saturating MMPS 5 100 1
SEVIR MMPS 1 100 1
@@
We used the Ensemble Score Filter (EnSF) as described in (Liang et al., 2025) with $\epsilon_{\alpha}=0.05$ and 1000 Euler steps. We note that the EnSF implementation includes an inflation step not mentioned in the original article, which we found crucial to prevent mode collapse. After each assimlation cycle, the particles are updated as
@@
We trained a conditional stochastic interpolant model for probabilistic forecasting (Chen et al., 2024), as used in FlowDAS, to emulate the forward dynamics of SQG at $64\times 64$ resolution. Our model conditions on the past six states for roll-out, i.e., generates samples from $p({\bm{x}}_{n}|{\bm{x}}_{n-1},\ldots,{\bm{x}}_{n-6})$. For the drift model $\bm{b}$}, we used a U-Net, similar to the architecture used for the interpolant in DAISI, with $64$ channels, and consisting of three hierarchical levels with channel multipliers $(1,2,2)$. Each level uses group-normalized residual blocks with 8 groups, and time embeddings are provided through learned sinusoidal features of dimension 32. Linear self-attention is applied at every resolution, together with a full multi-head self-attention block applied at the bottleneck between the encoder and decoder. The model is trained using AdamW with learning rate $2\times 10^{-4}$, cosine learning-rate decay, gradient-norm clipping, batch size 32, and dataset normalization with standard deviation $\sigma_{\text{train}}=2660$. Training is performed for 4096 epochs on 3-hourly SQG data.
For sampling, we use the parameters in Table 8, which are based on (Chen et al., 2025) and finetuned for the different cases.
@@
LETKF is a state-of-the-art DA method commonly used in the geosciences (Wang et al., 2021). The hyperparameters used for LETKF are shown in Table 9. These are roughly based on the ones in (Liang et al., 2025) but finetuned for the different cases. For the high-dimensional case, no parameter search is done due to computational cost.
@@
Non-stationary $1000\text{\,}\mathrm{km}$ 0.2
SEVIR $3\text{\,}\mathrm{p}\mathrm{x}$ 0.0